The cylinder has a density of 21186 kg m³ and hight of 39.17 mm. Calculate the cross sectional area of the cylinder


1
Expert's answer
2021-03-23T11:12:29-0400

Mass of the cylinder:


"m=\rho V."

Its volume:


"V=Ah,"

where A - the cross-cectional area. Thus,


"m=\rho Ah, \\\space\\
A=\frac{m}{\rho h}=\frac{m}{21186·0.03917}=\frac{m}{829.86}\text{ (m}^2)."